For high -pass active filter with corner frequency 8 kHz, then R and C are: 


Select one : 

a. 2.5kΩ ,7.9557nF 

b. 2 kΩ ,7.9557nF 

c. 5kΩ ,7.9557nF 

d. 10kΩ ,7.9557nF

e. 1kΩ ,7.9557nF


1
Expert's answer
2021-06-15T13:37:02-0400

"\\\nu=\frac{1}{2\pi RC}\to\frac{1}{2\cdot3.14\cdot2500\cdot7.9557\cdot10^{-9}}\approx8000\ (Hz)" . So,


a. 2.5kΩ ,7.9557nF. Answer